### v1

@@ 5.2 Vulnerability Classification
Following the code generation, we executed ESBMC on each file to classify them. Let us denote the set of all the generated C samples by $\Sigma$, such that $\Sigma=\{c_{1},c_{2},\ldots,c_{265,000}\}$, where each $c_{i}$ represents an individual sample. As shown in Table 3, we can group all the samples into four distinct categories. These categories are mutually exclusive, meaning a single sample cannot belong to more than one category.
Table 3: Summary of Vulnerabilities with Detailed Descriptions.
Category Description
$\mathcal{VS}$: Verification Success The set of samples for which the verification process completed successfully with no vulnerabilities detected.
$\mathcal{VU}$: Verification Unknown        (Timeout) The set of samples for which the verification process was not completed within the provided time frame; no vulnerabilities found before timeout.
$\mathcal{ER}$: Error The set of samples for which the verification status resulted in an error.
A significant change has been made compared to the classification in [9]. The previous paper only included vulnerabilities if a program’s verification process was completely finished. Our new approach includes every instance where ESBMC found a counterexample. This decision was motivated by identifying as many vulnerabilities as possible and providing the most accurate ratio of vulnerable to non-vulnerable programs. Dismissing these intermediate detection results could lead to a misleading assessment, potentially underestimating the number of vulnerabilities produced by larger models. Thus, including all counterexamples ensures a more accurate representation of the true vulnerability landscape across different LLM outputs. For this reason, the category “TIMEOUT” ($\mathcal{TO}$) has been renamed to $\mathcal{VU}$: Verification Unknown. The category labeled as “ERROR” ($\mathcal{ER}$) encompasses all instances where the verification process faced errors or crashes in the core ESBMC module, GOTO converter, SMT solver, or the clang compiler module. These samples are omitted from the classification because they cannot be handled using the latest ESBMC module. “Verification failed” ($\mathcal{VF}$) represents the main focus of our interest. This set of samples, where the verification status was unsuccessful, had vulnerabilities identified by ESBMC through counterexamples. We divided $\mathcal{VF}$ into $6$ main categories, and $19$ subcategories. Note that in the JSON file that contains the vulnerability classification, the exact type of vulnerability is always indicated as provided by the ESBMC module, as shown in Appendix Fig. 11. The categories and subcategories, along with the precise distribution of vulnerabilities across the entire dataset, are detailed in Section 6, Table 7.
@@ 5.2.1 ESBMC Parameter Selection
Model-checking tools like ESBMC offer adjustable parameters that can optimize performance based on program complexity. Default parameters such as those used in competitions like SV-COMP, may not be suitable for all software types, potentially leading to fewer detected vulnerabilities. We conducted a detailed analysis to understand how various settings impact verification outcomes, particularly for programs generated by LLMs. The different switches are explained in Appendix Table 16.
Given our expectation that it generates more complex code, we have selected $1,000$ samples generated by GPT-4, serving as the basis for selecting the ESBMC parameters for the entire dataset. For these samples, we tested multiple parameter configurations of ESBMC to determine which settings yielded the best results regarding runtime efficiency and vulnerability detection. We focus on two objectives. Firstly, to minimize verification unknown outcomes ($\mathcal{VU}$) through the $t$ parameter and preferably completing the verification process; and secondly, to identify as many vulnerabilities as possible. Table 4 illustrates the verification outcomes of the 1,000 samples, demonstrating how various combinations of unwind (u) and time (t), alongside the utilization of k-induction, incremental-bmc, or falsification techniques, impact the results. Our analysis revealed that merely increasing the unwind parameter $u$ while keeping a short timeout (e.g., 1 second) often leads to timeouts. For example, setting the unwind to $10$ with a 1-second timeout resulted in most samples (684) falling into $\mathcal{VU}$. This limitation reflects the capability of the underlying architecture; in our case, the AMD Ryzen Threadripper PRO 3995WX with 32 CPU cores. A more powerful system could potentially handle more computations within the same timeframe.
Table 4: Results of classification for varying parameters based on a dataset of 1000 samples generated by GPT-4.
@@
u time k-ind bmc fls Running time (m:s) $\boldsymbol{|\phi|}$ $\boldsymbol{\mathcal{VS}}$ $\boldsymbol{\mathcal{VF}}$ $\boldsymbol{\mathcal{VU}}$ $\boldsymbol{\mathcal{ER}}$
300 1698:53 1678 471 491 25 13
2 1000 1418:03 1638 505 407 70 18
3 1000 2100:36 1620 495 390 94 21
100 653:05 1583 486 468 33 13
2 100 224:25 1580 496 393 96 15
1 1000 419:45 1529 538 428 21 13
30 216:28 1513 494 448 45 13
30 216:20 1511 494 448 45 13
30 232:36 1511 494 448 45 13
2 30 99:05 1500 486 371 129 14
1 100 79:09 1465 536 421 30 13
10 84:11 1432 500 430 57 13
3 100 344:01 1408 478 350 158 14
1 10 21:47 1351 527 399 61 13
2 10 47:48 1272 469 330 187 14
3 10 62:14 951 433 272 281 14
1 13:23 941 474 336 177 13
1 13:39 938 475 335 177 13
1 13:34 936 476 334 177 13
1 1 7:41 911 487 323 177 13
2 1 10:30 559 404 205 377 14
10 1 14:14 158 224 79 684 13
10 152:41 69 75 25 887 13
@@
✓: Enabled; ✗: Not set; $|\phi|$: Number of Vulnerabilities; $\boldsymbol{\mathcal{VS}}$: Verification Success;
$\boldsymbol{\mathcal{VF}}$: Verification Failed: vulnerabilities detected; $\boldsymbol{\mathcal{VU}}$: Verification Unknown (Timeout) $\boldsymbol{\mathcal{ER}}$: Error: Verification process resulted in an error; k-ind: k-induction;
bmc: incremental-bmc; fls: falsification technique; u: unwind
A larger unwind parameter increases the detection of vulnerabilities in loops if adequate processing time is given. K-induction improves the detection rate for larger unwind settings, but the default k-bound is 1 unless adjusted. After extensive testing, we extended the timeout to 500 seconds and allowed unlimited k-steps, transitioning from bounded to unbounded model checking. This adjustment ensures that if the verification is completed within this time frame, we either identify a counterexample or confirm the absence of the examined vulnerabilities. This represents a significant change from our previous methodology in creating the FormAI-v1 dataset. We used the following ESBMC switches during our experiments, as depicted in Figure 7.
@@
Note that --overflow, --memory-leak-check, and --multi-property are the same for the entire Table 4 experiments and were not changed. Using these parameters on our $1000$ sample set, 416 files were deemed non-vulnerable, while 519 files were vulnerable. Among these 519 files, a total of 2116 unique vulnerabilities were detected. Considering the classification of $256,000$ programs, the worst-case scenario is that every program from FormAI-v2 would utilize its allocated time, resulting in 500 seconds dedicated to verifying each sample. Using 32 CPU threads, the entire verification process on our experimental setup would take approximately 46,3 days in this worst-case scenario, calculated as $256,000\times 500/60/60/24/32$.
It is important to notice that while the best-performing setting from Table 4 had a much lower time setting than the second best, more files utilized the maximum time frame, thus making the verification process longer.
6 Discussion
The following section summarizes our main results. First, we examine statistics on the entire dataset regarding overall verification results and vulnerability types. This is followed by evaluating each LLM and comparing their cyclomatic complexity, keyword frequency and secure coding capabilities. In the original FormAI dataset, we only created $112.000$ compilable C samples using GPT-3.5. Furthermore, the complexity of each program was not measured. This research closes this gap by comparing eight state-of-the-art LLMs and providing a vulnerability-labelled dataset to the research community, comprising $256.000$ C programs.
6.1 Evaluation of the FormAI-v2 Dataset
We have examined over $21,994,613$ lines of C code, with an average of $83.70$ lines per sample. In total, we performed the verification process on $256,000$ C program files, and our results for the entire dataset are shown in Table 5. The TOP 10 violations throughout the entire dataset are presented in Table 6.
Table 5: Summary of Verification Outcomes.
Category Count (%)
Total Files 265,000 100.00
Verification Success 23,036 8.69
Verification Unknown (Timeout) 66,899 25.24
Verification Failed 168,208 63.47
Error 6,857 2.59
Table 6: Total Violations Across All Categories.
Rank Category Violation Type Count Percentage
1 $\mathcal{DF}$ Dereference failure: NULL pointer 285,505 41.73%
2 $\mathcal{BO}$ Buffer overflow on scanf 181,122 26.47%
3 $\mathcal{DF}$ Dereference failure: invalid pointer 61,022 8.92%
4 $\mathcal{DF}$ Dereference failure: forgotten memory 21,222 3.10%
5 $\mathcal{DF}$ Dereference failure: array bounds violated 20,200 2.95%
6 $\mathcal{DF}$ Array bounds violated: upper bound 18,895 2.76%
7 $\mathcal{DF}$ Array bounds violated: lower bound 17,275 2.52%
8 $\mathcal{AO}$ Arithmetic overflow on sub 15,089 2.21%
9 $\mathcal{AO}$ Arithmetic overflow on add 12,552 1.83%
10 $\mathcal{AO}$ Arithmetic overflow on mul 10,088 1.47%
During the $500$-second verification time-frame, ESBMC identified $168,208$ unique programs with vulnerabilities. In contrast, only $23,036$ programs, representing $8.69\%$, were verified as secure. Expanding computational resources may increase the number of programs uncovered from $\mathcal{VU}$, thereby potentially extending the $\mathcal{VF}$ category. These given results provide an even better lower bound compared to [9], on what percentage of LLM-generated code is vulnerable. The big picture is worse than simply saying $63.47\%$, as one file can contain more than one vulnerability. The total number of property violations detected by ESBMC is $684,227$. A breakdown of the distribution of these different types of vulnerabilities is shown in Table 7.
Table 7: Detailed Categorisation of Vulnerabilities as Detected by ESBMC 7.5.
Category Description Count Percentage
$\mathcal{DF}$ Dereference failures:
   - NULL pointer 285505 41.73%
   - Invalid pointer 61022 8.92%
   - Forgotten memory 21222 3.10%
   - Array bounds violated 20200 2.95%
   - Invalidated dynamic object 2904 0.42%
   - Access to object out of bounds 1898 0.28%
   - Accessed expired variable pointer 1091 0.16%
   - Write access to string constant 751 0.11%
   - Of non-dynamic memory 251 0.04%
   - Object accessed with incompatible base type 270 0.04%
   - Oversized field offset 141 0.02%
   - Data object accessed with code type 11 0.00%
$\mathcal{AO}$ Arithmetic overflows:
   - On sub 15089 2.21%
   - On add 12552 1.83%
   - On mul 10088 1.47%
   - IEEE mul 7545 1.10%
   - IEEE div 2755 0.40%
   - IEEE add 1729 0.25%
   - IEEE sub 1464 0.21%
   - On div 691 0.10%
   - On shl 669 0.10%
   - On modulus 220 0.03%
   - On neg 105 0.02%
$\mathcal{BO}$ Buffer overflows:
   - On scanf 181122 26.47%
   - On fscanf 6593 0.96%
   - On sscanf 2810 0.41%
$\mathcal{ABV}$ Array bounds violations:
   - Upper bound 18895 2.76%
   - Lower bound 17275 2.52%
   - VLA array size in bytes overflows address space size 4082 0.60%
$\mathcal{DBZ}$ Division by zero 3592 0.52%
$\mathcal{MV}$ Miscellaneous Vulnerabilities:
   - The pointer to a file object must be a valid argument 1125 0.16%
   - Invalid Function argument issues 339 0.05%
   - Same object violation 137 0.02%
   - Operand of free must have zero pointer offset 84 0.01%
The most common type of vulnerability is related to “Dereference failures” accounting for $57.77\%$ of the cases, predominantly due to NULL pointer issues. This category includes a variety of pointer-related issues such as invalid pointers, forgotten memory, and array bounds violations, among others. “Buffer overflows”, mainly triggered by the scanf function, comprise a significant 27.84% of the vulnerabilities. This highlights common issues in handling buffer sizes and input functions. “Arithmetic overflows” are also notable, covering various operations like subtraction, addition, multiplication, and division, indicating frequent issues in handling numeric calculations without adequate checks. The table further lists “Array bounds violations” and “Division by zero” as common issues, illustrating challenges in correctly managing arrays and arithmetic operations. A smaller portion of the table covers ”Miscellaneous Vulnerabilities,” which includes a variety of less frequent but notable issues such as invalid file object pointers and operand violations in memory deallocation. Overall, the data emphasizes the need for robust handling of pointers, buffers, and numeric operations within the source code to mitigate the risk of vulnerabilities.
6.2 CWE Classification
Next, we link the vulnerabilities to Common Weakness Enumeration (CWE) identifiers by manually assigning the appropriate CWEs after reviewing source codes associated with each vulnerability group. The multifaceted nature of software flaws often results in a single vulnerability associated with multiple CWE identifiers. Table LABEL:tab:vulnerability shows the categorization of the most prevalent vulnerabilities and the associated 39 unique CWEs we identified across these categories. The “Miscellaneous Vulnerabilities” ($\mathcal{MV}$) category includes Assertion failure, Same object violation, Operand of free must have zero pointer offset, function call: not enough arguments, and several types of dereference failure issues. For these categories we did not individually assign CWE identifiers, as in general they constitute a very small percentage of the dataset. The small sample size also prevents machine learning algorithms to efficiently recognise and learn these patterns. While in the .json file they are still marked as vulnerable and count for statistical purposes, we aggregated these samples under the other category so they can be conveniently excluded from the training data.
In total, $42$ unique CWE were identified in the dataset. From MITRE’s Top 25 Most Dangerous Software Weaknesses for 2023 list, six is present in our list as shown in Table 8.
Table 8: CWEs from 2023’s MITRE Top 25.
Rank CWE Description
1 CWE-787: Out-of-bounds Write
4 CWE-416: Use After Free
6 CWE-20: Improper Input Validation
7 CWE-125: Out-of-bounds Read
12 CWE-476: NULL Pointer Dereference
14 CWE-190: Integer Overflow or Wraparound
The remaining CWEs in the top 25 list are related to web vulnerabilities like SQL injection, XSS, and authentication, which are irrelevant to our C language samples. It is vital to emphasize that, in our situation, classifying the C programs based on CWE identifiers is not practical, contrary to what is done for other databases like Juliet. As shown in Table 1, most datasets contain only one vulnerability per sample. In the datasets ReVeal, BigVul, and Diversevul, a function is vulnerable if the vulnerability-fixing commit changes it, while in Juliet, a single vulnerability is introduced for each program. In FormAI, a single file often contains multiple vulnerabilities. As noted, a single vulnerability can be associated with multiple CWEs. Additionally, multiple CWEs can be required as a prerequisite for a vulnerability to manifest. As an example, “CWE-120: Buffer Copy without Checking Size of Input (Classic Buffer Overflow)”, can happen as a result of “CWE-676: Use of Potentially Dangerous Function”, which can be the scanf function. If this is combined with “CWE-20: Improper Input Validation”, it can result in “CWE-787: Out-of-bounds Write”.
Table 9: Detailed Categorisation of Vulnerabilities as Detected by ESBMC 7.5.
Description CWE Associated CWEs
Dereference failures:
 - NULL pointer CWE-476 CWE-690, CWE-391
 - Invalid pointer CWE-822 CWE-119, CWE-787, CWE-822
 - Forgotten memory CWE-825 CWE-401, CWE-404, CWE-459
 - Array bounds violated CWE-125 CWE-119, CWE-787
 - Invalidated dynamic object CWE-824 CWE-416, CWE-415
 - Access to object out of bounds CWE-125 CWE-119, CWE-787
 - Accessed expired variable pointer CWE-416 CWE-825
 - Write access to string constant CWE-843 CWE-758
 - Of non-dynamic memory CWE-590 CWE-415, CWE-415, CWE-762
 - Object accessed with incompatible base type CWE-843 CWE-119
 - Oversized field offset CWE-787 CWE-119, CWE-125, CWE-823
 - Data object accessed with code type CWE-843 CWE-686, CWE-704
Arithmetic overflows:
 - On sub CWE-191 CWE-20, CWE-190, CWE-192
 - On add CWE-190 CWE-20, CWE-191, CWE-192
 - On mul CWE-190 CWE-20, CWE-191, CWE-192
 - Floating-point ieee_mul CWE-190 CWE-681
 - Floating-point ieee_div CWE-682 CWE-369, CWE-681
 - Floating-point ieee_add CWE-190 CWE-681
 - Floating-point ieee_sub CWE-190 CWE-681
 - On div CWE-190 CWE-20, CWE-369
 - On shl CWE-190 CWE-192
 - On modulus CWE-190 CWE-20, CWE-191
 - On neg CWE-191 CWE-190, CWE-192
Buffer overflows:
 - On scanf CWE-120 {CWE-20, CWE-121, CWE-122
 - On fscanf CWE-120 CWE-129, CWE-131, CWE-628
 - On sscanf CWE-120 CWE-676, CWE-680, CWE-787}
Array bounds violations:
 - lower bound CWE-129 {CWE-119, CWE-125, CWE-129
 - upper bound CWE-788 CWE-131, CWE-193, CWE-787}
 - VLA array size in bytes overflows CWE-190 CWE-131, CWE-680
Division by zero CWE-369 CWE-691
Miscellaneous Vulnerabilities:
 - The pointer to a file must be valid CWE-476 CWE-690, CWE-459
 - Same object violation CWE-628 CWE-843, CWE-668
 - Operand of free must have zero pointer offset CWE-761 CWE-415, CWE-590
Labelling the vulnerable function name, line number, and vulnerability type identified by the ESBMC module provides granular information that can benefit machine learning training. This level of detail can allow models to discern patterns and correlations with higher precision, thereby improving vulnerability prediction and detection capabilities. Since our programs exhibit numerous vulnerabilities, including multiple occurrences of the same type, categorizing each solely into one CWE group, as seen with Juliet, would be sub-optimal for training purposes. This method fails to communicate crucial details about the vulnerabilities. For instance, both ”Arithmetic overflow on add” and ”Arithmetic overflow on div” are assigned the same primary CWE, they manifest differently in the source code. Therefore, merely labelling them with CWEs does not offer the same level of granularity and makes the dataset less suitable for ML. While other datasets focus more on CWEs related to vulnerabilities that could potentially be exploited, ESBMC also detects issues related to software safety.
6.3 LLMs comparison
6.3.1 Keyword Frequency
We employ a token-based keyword-counting mechanism to extract the cardinality of the $32$ C keywords from each LLM-generated subset, as shown in Figure 8 and 9. The frequency is normalized to show the occurrence of keywords per million lines of code for each model.
Figure 8: 32 C keyword distribution - PART I.
It is also important to note that while the distribution is similar, the slight fluctuations indicate that the different models handle statements, expressions, and variables differently. This observation is supported by examining the cyclomatic complexity statistics of each model.
Figure 9: 32 C keyword distribution - PART II.
6.4 Cyclomatic Complexity and General Statistics
To comprehend the meaning and limitations of cyclomatic complexity, it’s essential to delve into the industry-standard guidelines established by Microsoft and NIST [109, 110]. Simply put, “higher numbers are bad and lower numbers are good”. Cyclomatic complexity serves as a metric for assessing the difficulty involved in testing, maintaining, or troubleshooting code, as well as predicting the likelihood of errors. It is calculated by counting decision points in the source code. NIST defines it as“the amount of decision logic in a source code function” and sets the recommended maximum as $10$ [110]. They note that a complexity threshold above 10 is only acceptable for bigger projects where experienced engineers can manage the extra testing needed. Since the programs are small standalone units, a cyclomatic complexity number surpassing 10 is deemed unacceptable, and even the suggested threshold may be considered excessive.
Table 10: Statistics of Cyclomatic Complexity, ordered by Average CC.

### v2

@@ 5.3 Vulnerability Classification
After code generation and elimination, the next step was classifying our dataset using ESBMC. Compared to the classification in [9], a significant change has been made. The original work used bounded model checking (BMC) with a bound set to $k=1$. For example, if a property violation occurs at level $k=2$, then $BMC_{\Phi}(1)$ will not detect the vulnerability and it simply returns Verification Success, giving a false impression. As a result, in the FormAI-v1 dataset, numerous samples were previously classified as “NON-VULNERABLE up to bound $k$. We have transitioned from bounded to unbounded model checking to capture more vulnerabilities or prove their absence for each sample. This approach incrementally unwinds the program until a bug is found or the completeness threshold is reached, meaning all possible terminating states have been explored. Incremental BMC ensures that smaller problems are solved sequentially, avoiding the need to guess an upper bound for verification.
Applying these settings, we have successfully identified more vulnerabilities in the programs. Consequently, if the verification process is completed successfully, we can conclude that the program has no violated properties (that can be detected by the currently used ESBMC version). While this approach requires significantly more computational power, it has proven effective in revealing more vulnerabilities or proving their absence, as we will demonstrate in Section 6.
@@ 5.3.1 ESBMC Parameter Selection
Model-checking tools like ESBMC provide various parameters, and the identified vulnerabilities may differ depending on the parameters chosen. Default parameters, such as those used in competitions like SV-COMP, may not be suitable for all software types, potentially leading to fewer detected vulnerabilities. This naturally leads to the question: which options should we use? Should we use the k-induction switch with a large time limit, such as 100 seconds, or should we opt for the bmc switch? These questions are not straightforward. To address them, we conducted a detailed analysis to understand how different settings impact verification outcomes. We have randomly selected 1 00010001\,0001 000 samples from the dataset, serving as the basis for selecting the ESBMC parameters for the entire dataset. By experimenting with different switches and time frames, we were able to select the options that best met our needs.
For these samples, we tested multiple parameter configurations of ESBMC to determine which settings yielded the best results regarding runtime efficiency and vulnerability detection. We focus on two objectives. Firstly, to minimize verification unknown outcomes ($\mathcal{VU}$) through the $t$ (time) parameter and preferably completing the verification process; and secondly, to identify as many vulnerabilities as possible. Table 7 illustrates the verification outcomes of the 1,000 samples, demonstrating how various combinations of unwind (u) and time (t), alongside the utilization of k-induction, incremental BMC, or falsification techniques, impact the results.
Table 7: Classification Results for the 1000-Sample Dataset With Varying Parameters.
@@
[…]
x 300 x x 1698:53 1 67816781\,6781 678 $471$ $491$ $25$ $13$
2 1000 x x x 1418:03 1 63816381\,6381 638 $505$ $407$ $70$ $18$
3 1000 x x x 2100:36 1 62016201\,6201 620 $495$ $390$ $94$ $21$
x 100 x x 653:05 1 58315831\,5831 583 $486$ $468$ $33$ $13$
2 100 x x x 224:25 1 58015801\,5801 580 $496$ $393$ $96$ $15$
1 1000 x x x 419:45 1 52915291\,5291 529 $538$ $428$ $21$ $13$
x 30 x x 216:28 1 51315131\,5131 513 $494$ $448$ $45$ $13$
x 30 x x 216:20 1 51115111\,5111 511 $494$ $448$ $45$ $13$
x 30 x x 232:36 1 51115111\,5111 511 $494$ $448$ $45$ $13$
2 30 x x x 99:05 1 50015001\,5001 500 $486$ $371$ $129$ $14$
1 100 x x x 79:09 1 46514651\,4651 465 $536$ $421$ $30$ $13$
x 10 x x 84:11 1 43214321\,4321 432 $500$ $430$ $57$ $13$
3 100 x x x 344:01 1 40814081\,4081 408 $478$ $350$ $158$ $14$
1 10 x x x 21:47 1 35113511\,3511 351 $527$ $399$ $61$ $13$
2 10 x x x 47:48 1 27212721\,2721 272 $469$ $330$ $187$ $14$
3 10 x x x 62:14 $951$ $433$ $272$ $281$ $14$
x 1 x x 13:23 $941$ $474$ $336$ $177$ $13$
x 1 x x 13:39 $938$ $475$ $335$ $177$ $13$
x 1 x x 13:34 $936$ $476$ $334$ $177$ $13$
1 1 x x x 7:41 $911$ $487$ $323$ $177$ $13$
2 1 x x x 10:30 $559$ $404$ $205$ $377$ $14$
10 1 x x x 14:14 $158$ $224$ $79$ $684$ $13$
x 10 x x x 152:41 $69$ $75$ $25$ $887$ $13$
@@
✓: Enabled; x: Not set; $|\phi|$: Number of Vulnerabilities detected; k-ind: k-induction;
bmc: incremental-bmc; fls: falsification technique; u: unwind; (Runtime in (m:s))
In this context, “unwind” refers to the number of iterations for which we should unroll the loops. For example, u = 1 means that we unroll a loop for only one iteration, while u = $\infty$ indicates no limit on the number of iterations for loop unrolling. Our analysis revealed that merely increasing the unwind parameter $u$ while keeping a short timeout (e.g., 1 second) often leads to timeouts. For example, setting the unwind to $10$ with a 1-second timeout resulted in most samples (684) falling into $\mathcal{VU}$. A larger unwind parameter enhances the detection of vulnerabilities in loops, provided there is sufficient processing time. We can also observe that the k-induction switch increases the number of detected vulnerabilities, $|\phi|$, as the allotted time increases. Therefore, the best approach for us is to set the timeout as high as possible with k-induction enabled. In our architecture, we set the timeout to 500 seconds and allowed unlimited k-steps, transitioning from bounded to unbounded model checking. This adjustment ensures that if the verification is completed within this time frame, we either identify a counterexample or confirm the absence of the examined vulnerabilities. For our dataset classification, we used a machine with 192 CPUs and 1.5 TB of memory, which allowed us to set a time frame of 500. A larger time frame is not feasible in our test environment, as concurrently verifying 192 C programs would exceed the 1.5 TB memory capacity. Based on our experiments, we used the following ESBMC switches during our experiments, as depicted in Figure 6.
@@
Note, that the --overflow, --memory-leak-check, and --multi-property switches are used to identify the maximum number of potential vulnerabilities. These switches do not affect the running time. Using these parameters on our $1000$ sample set, 416 files were deemed non-vulnerable, while 519 files were vulnerable. Among these 519 files, a total of 2116 unique vulnerabilities were detected. Considering the classification of 331 000331000331\,000331 000 programs, the worst-case scenario is that every program from FormAI-v2 would utilize its allocated time, resulting in 500 seconds dedicated to verifying each sample. Using 192 CPU threads, the entire verification process on our experimental setup would take approximately $9,97$ days in this worst-case scenario, calculated as 331 000×500/60/60/24/192331000500606024192331\,000$\times$ 500/60/60/24/192331 000 × 500 / 60 / 60 / 24 / 192.
6 Verification Results
In this section, we summarize our key results, beginning with an analysis of statistics for the entire dataset, focusing on overall verification outcomes and vulnerability types. It is important to note that the analysis is based on $310,531$ programs, as all clones up to Type 3-2 have been excluded from the initial $331,000$. We then evaluate each LLM, comparing the complexity of the code they generate and their secure coding capabilities. This is followed by evaluating each LLM and comparing the complexity of the code they generate and their secure coding capabilities.
In the original FormAI dataset, only $112,000$ compilable C samples were created using GPT-3.5-turbo. Furthermore, the complexity of each program was not measured. This research closes this gap by comparing nine state-of-the-art LLMs and providing a vulnerability-labelled dataset to the research community. We have examined 26 633 1562663315626\,633\,15626 633 156 lines of C code, with an average of $85.77$ lines per sample. In total, we performed the verification process on 310 531310531310\,531310 531 C program files, and our results for the entire dataset are shown in Table 8. The TOP 10 violations throughout the entire dataset are presented in Table 9. Table 10 provides a breakdown of the distribution for each of the top five main categories of vulnerabilities.
Table 8: Overview of Statistics and Verification Results for Each LLM.